The current in a long solenoid of radius 3 cm and 471 turns per meter is varied with time at a rate of 2 A/s. A circular loopPLEASE SHOW WORK!
Answer #1

Now Mutual Inductance is given by M = Mo N, N₂A Azarea of inner e solenoid M= (44*10-2)(4716) a(0.03m)2 N = 471 M = 16.71 * 1